OpenStudy (anonymous):

what is the cubed root of -1

OpenStudy (anonymous):

\[x^3 + 1 = 0 \] \[(x + 1 )(x^2 + 1 - x)=0\]

OpenStudy (anonymous):

\(x= - 1\) is one of the roots and also the only real root...

OpenStudy (anonymous):

For the complex part we need to use quadratic formula.

OpenStudy (chaise):

\[\sqrt[3]{-1}=-1^(1/3) = -1\]

OpenStudy (anonymous):

\[x^2+ 1 -x = 0\] \[x = \frac{1 \pm i\sqrt3}{2}\]

OpenStudy (anonymous):

Now you have three roots \( -1 , \frac{1 + i\sqrt3}{2} , \frac{1 - i\sqrt3}{2}\)